# Qiumin Zhu, Zhen Sun, Songpengcheng Xia, Guoqing Liu, Kehui Ma, Ling Pei†∗ and Zheng Gong§∗

### v1

@@ I Introduction
Therefore, the other challenge concerns the definition of traversability cost labels. Although a variety of works propose learning-based costmap approaches, they predict the cost ignoring the nuance of the robot’s interactions with different terrains. Some simply divide the ground into traversable and untraversable whether the vehicle can reach or not [8], while others represent slip as the velocity error [9]. To capture roughness and bumpiness that the vehicle experiences, proprioceptive sensors like Inertial Measurement Units (IMUs), are useful for sensing these characteristics. It is easy for IMUs to build on the ground vehicle and capture the state [10, 11]. The signals reflect the vehicle’s vibrations, movements, and changes in feelings as traversability costs. We take the properties of IMU data and robotic risk into consideration, handling the IMU linear acceleration in the z-axis to generate continuous traversability values as learning targets.
@@
II RELATED WORK
To estimate traversability, hybrid methods encode both geometric and semantic information to build a traversability map representing the vehicle’s surroundings. [12, 13] compute continuous geometric traversability scores, assign discrete scores to various semantics and calculate traversability cost with the sum of these two scores.
Recently, many works have focused on costmap learning for autonomous navigation in challenging off-road environments. Fan et al. [14] learn a traversability risk-aware costmap through a CNN with LiDAR point clouds as inputs and geometric cost as ground truth labels. Cai et al. [15] learn a speed distribution map from a semantic input and convert the map into a costmap with a conditional value at risk (CVaR). Seo et al. [8] leverage Positive-Unlabeled (PU) learning method and a 2D normalizing flow to learn a binary traversability map considering wheel-contact points as traversable. Although these approaches predict costmaps using visual and geometric characteristics captured by cameras and LiDARs, the interactions between the robot and the ground surface are ignored in the process of cost prediction. To learn a costmap based on the interaction, Frey et al. [9] use the discrepancy between the robot’s current linear velocity and reference linear velocity as labels to estimate dense traversability from RGB images.
Besides the robot’s linear velocity, other interactions can be sensed by IMU. Sathyamoorthy et al. [16] generate the cost labels by applying Principal Component Analysis (PCA) to reduce the dimensions of the 6-dimensional IMU data to two principal components. Waibel et al. [17] combine normalized angular velocity in x and y and linear acceleration in z to obtain the real IMU cost. Seo et al. [18] project the magnitude of z-acceleration from IMU to contact points as ground-truth traversability. Some works process the IMU data at frequency domain. Yao et al. [19] analyze the average amplitude spectrum of the x-axis and y-axis angular velocity and the z-axis linear acceleration by Fast Fourier Transform (FFT) to calculate traversable costs. Castro et al. [20] describe the traversability as properties of the terrain with the bandpower of the z-axis IMU linear acceleration. We use colored point cloud data to populate continuous costmaps, training with risk-aware IMU traversability cost labels and demonstrate our approach in complex and challenging off-road terrains on the autonomous robot platform.
III METHODS
III-A Overview
@@
III-B Traversability Labels Generation
@@
III-C 3D Environment and Robot Motion Preprocessing
@@ III-D Costmap Prediction
IV EXPERIMENT AND RESULTS
IV-A Simulation Setup
@@ IV-C Costmap Evaluation
TABLE I: Results of Costmap Comparisons
@@
Table I shows the comparison of costmap prediction performance in two scenarios. Our method has better performance on accuracy and MSE. Fig. 5 indicates that our method performs similarly to TNS on AUC, both superior to HDIF. TNS divides the semantic classes into traversable and untraversable terrains in advance, hence it has a little advantage over ours. The result of the costmap comparisons verifies that our definition of traversability cost is reasonable and our costmap prediction is accurate.
IV-D Navigation Evaluation
@@
TABLE II: Results of Navigation Comparisons
@@
Table II shows the comparison of navigation performance in two scenarios. Our method outperforms the other methods in terms of all four metrics. HDIF takes shorter trajectories in the forest scene since it travels directly towards the goal without considering the obstacles and its successful trajectories for evaluation are less than ours. A higher success rate of our method shows that it is effective even in different and complex situations while the other methods have failed. The shorter normalized trajectory length of our method indicates that it is more efficient and precise by minimizing unnecessary movements and simultaneously avoiding collision. The less time cost and higher stability in the condition of the same preset speed demonstrate that our method ensures a faster but safer and more stable navigation. The trajectories of the navigation experiment are illustrated in Fig. 6.
IV-E Further Analysis
We use angular velocity as the inputs and an LSTM in the network. To validate whether they improve the performance of our network, we make an ablation study. We use the same dataset to train three networks: ours, ours without angular velocity $\omega$ and ours without LSTM. We compare the three best learned models in the validation set and test set. Table III shows that angular velocity and LSTM both have improvements in the loss of the network.
TABLE III: Network Loss Comparisons

### v2

@@ 1 Introduction
Therefore, the other challenge concerns the definition of traversability cost labels. Although a variety of works propose learning-based costmap approaches, they predict the cost ignoring the nuance of the robot’s interactions with different terrains. Some simply divide the ground into traversable and untraversable whether the vehicle can reach or not [8], while others represent slip as the velocity error [9]. To capture roughness and bumpiness that the vehicle experiences, proprioceptive sensors like Inertial Measurement Units (IMUs), are useful for sensing these characteristics. It is easy for IMUs to build on the ground vehicle and capture the state [10]. The signals reflect the vehicle’s vibrations, movements, and changes in feelings as traversability costs. We take the properties of IMU data and robotic risk into consideration, handling the IMU linear acceleration in the z-axis to generate continuous traversability values as learning targets.
@@
2 Methods
2.1 Overview
@@
2.2 Traversability Labels Generation
@@
2.3 3D Environment and Robot Motion Preprocessing
@@ 2.4 Costmap Prediction
3 Experiment and Results
3.1 Simulation Setup
@@ 3.3 Costmap Evaluation
Table 1: Results of Costmap Comparisons
@@
Table 1 shows the comparison of costmap prediction performance in two scenarios. Our method has better performance on accuracy and MSE. Fig. 5 indicates that our method performs similarly to TNS on AUC, both superior to HDIF. TNS divides the semantic classes into traversable and untraversable terrains in advance, hence it has a little advantage over ours. The result of the costmap comparisons verifies that our definition of traversability cost is reasonable and our costmap prediction is accurate.
3.4 Navigation Evaluation
@@
Table 2: Results of Navigation Comparisons
@@
Table 2 shows the comparison of navigation performance in two scenarios. Our method outperforms the other methods in terms of all four metrics. HDIF takes shorter trajectories in the forest scene since it travels directly towards the goal without considering the obstacles and its successful trajectories for evaluation are less than ours. A higher success rate of our method shows that it is effective even in different and complex situations while the other methods have failed. The shorter normalized trajectory length of our method indicates that it is more efficient and precise by minimizing unnecessary movements and simultaneously avoiding collision. The less time cost and higher stability in the condition of the same preset speed demonstrate that our method ensures a faster but safer and more stable navigation. The trajectories of the navigation experiment are illustrated in Fig. 6.
3.5 Further Analysis
We use angular velocity as the inputs and an LSTM in the network. To validate whether they improve the performance of our network, we make an ablation study. We use the same dataset to train three networks: ours, ours without angular velocity $\omega$ and ours without LSTM. We compare the three best learned models in the validation set and test set. Table 3 shows that angular velocity and LSTM both have improvements in the loss of the network.
Table 3: Network Loss Comparisons
@@ 4 Conclusions
Acknowledgement
This work was supported in part by the National Natural Science Foundation of China (NSFC) under Grant No.62273229 and smart city beidou spatial-temporal digital base construction and application industrialization (HCXBCY-2023-020).